OpenStudy (anonymous):

How do you do linear equations? ex. 10(3x + 13) + 8x - 15

OpenStudy (uri):

10*3x +10*13? @FallinLeaf

OpenStudy (anonymous):

I just started these and the teacher didn't explain them well. So I was wondering if someone could show me how.

OpenStudy (uri):

what is 10*3? @FallinLeaf

OpenStudy (anonymous):

Do you what me to multiply it or what. I a little confused.

OpenStudy (uri):

Multiply,yes.

OpenStudy (anonymous):

It is 30

OpenStudy (uri):

Okay,and what is 10*13

OpenStudy (anonymous):

130

OpenStudy (uri):

So,you now have 30x+130+8x-15,right? @FallinLeaf

OpenStudy (anonymous):

yeah, I think im starting to get it a little

OpenStudy (uri):

What is 30x+8x? @FallinLeaf

OpenStudy (anonymous):

I don't know what the x's mean.

OpenStudy (uri):

x+x =2x x*x=x^2

OpenStudy (uri):

Just add 30+8

OpenStudy (uri):

So,that gives you 38x.

OpenStudy (anonymous):

ok it would be 38 then.

OpenStudy (uri):

And then,what's 130-15?

OpenStudy (anonymous):

115

OpenStudy (uri):

So you get your answer. 38x+115.

OpenStudy (anonymous):

Thank you sooo much I was really getting frustrated. I was also wondering what do you do with the medals?

OpenStudy (skullpatrol):